OpenStudy (anonymous):

Can someone help me factor 2b^2+3b-2? It's been awhile sinece I last factored.

OpenStudy (anonymous):

(b+2)(2b−1)

OpenStudy (anonymous):

thank you!

OpenStudy (whpalmer4):

Yes, here's how to do it: \[2b^2+3b-2\] Multiply the coefficients of the first and last terms: 2*(-2) = -4 Now, factor -4 and look for a pair of factors which adds to 3: -1*4 = -4, and -1 + 4 = 3, so that's our pair. Rewrite the original polynomial, splitting the middle term with those factors: \[2b^2- 1b + 4b -2\]Group each pair of terms: \[(2b^2-1b) + (4b-2)\]Factor each group \[b(2b-1) + 2(2b-1)\]Notice that \((2b-1)\) is a common factor to both, and factor it out: \[(2b-1)(b+2)\] Check by multiplication: \[(2b-1)(b+2) = 2b(b+2) - 1(b+2) = 2b^2 + 4b -1b -2 = 2b^2 +3b -2 \checkmark\]